Exercise 18-19 Your answer is partially correct. Try again. On June 3, 2017, Sweet Company sold to Ann Mount merchandise having a sales price of $8,300 (cost $6,640) with terms of n/60, f.o.b. shipping point. Sweet estimates that merchandise with a sales value of $830 will be returned. An invoice totaling $100 was received by Mount on June 8 from Olympic Transport Service for the freight cost. Upon receipt of the goods, on June 8, Mount retuned to Sweet $300 of merchandise containing flaws. Sweet estimates the returned items are expected to be resold at a profit The freight on the returned merchandise was $24, paid by Sweet on June 8. On July 16, the compary received a check for the balance due from Mount Prepare journal entries for Sweet Companry to record all the events in June and July. (Credit account tities are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No entry for the account tities and enter 0 for the amounts.) Date Account Titles and Explanation Debit Credit Jun. 3, 2017Accounts Receivable Sales Revenue To recard sales) Cost of Goods Sold To record cost of goods sold) Jun. B, 2017Sales Returns and Allowances To record sales returrB) 17 Cost of Goods Sold To record cost of goods retuned) Jun. 8, 2017 ▼ | | Delivery Expense Cash To record the freight cost) Jul. 16, 2017Cash Accounts Receivable Click if you would like to Show Work for this question: Open Show Worik
Answer #1

Answer

  • All entries by you are correct, except for the June 8, 2017 one regarding the “Cost of Goods return”
  • The error is ONLY in the amount, which you calculated and filled as $ 176.
  • Calculation of correct amount:

--- Total cost of goods sold of original sale = $ 6,640
--- Sale value of original sale = $ 8,300

---Sale Value of Goods Returned = $ 300

---Cost of Goods Returned = (Cost of Goods Sold of original sale/Original sale value) x Sale value of Goods returned.

= (6640 / 8300) x 300 = $ 240

  • Correct figure = $ 240 to be filled (in place of $ 176)
  • OR, the same can also be calculated as follow:

--- Total cost of goods sold of original sale = $ 6,640
--- Sale value of original sale = $ 8,300
---% of cost of Goods Sold = 6640 / 8300 = 0.8 = 80%

---Sale value of goods returned = $ 300
---Cost of Goods returned = $ 300 x 80% = $ 240 = Answer.